# PRISM-DP: Spatial Pose-based Observations for Diffusion-Policies via Segmentation, Mesh Generation, and Pose Tracking

### v1

@@ 1 Introduction
Diffusion-based visuomotor policies generate robot motions by learning to denoise action-space trajectories conditioned on observations [1]. These observations are commonly sequences of raw RGB images, appealing due to its generality and ease of deployment, requiring just a camera. However, raw image sequences contain substantial amounts of irrelevant information, making it challenging for models to extract meaningful patterns without extensive computational resources, often necessitating hundreds of millions of parameters to fit demonstration data effectively.
@@
Recent work in learning-based 6D pose estimation and tracking, notably FoundationPose from Wen et al. [2], have introduced promising solutions for accurately estimating and tracking object poses directly from RGB-D video streams. In theory, these methods could enable compact pose-based diffusion policy training without external tracking setups. However, a significant limitation remains: FoundationPose and similar approaches depend on manually reconstructed or hand-crafted object meshes as input. Prior attempts to integrate FoundationPose into diffusion policy frameworks [3] have been constrained by the need for manual preprocessing of object meshes, hindering scalability in dynamic, open-set environments where object composition varies across tasks.
To address these challenges, we introduce Pose-tRacking via Image Segmentation and Mesh-generation Diffusion Policies (PRISM-DP), an approach that leverages segmentation, mesh generation, and pose estimation, and pose tracking models to enable compact diffusion policy learning directly from the spatial poses of task-relevant objects. More specifically, PRISM-DP applies a segmentation model to isolate task-relevant objects in input images, it uses the resulting segmentation masks to automatically create 3D triangulated meshes for all target objects using a mesh generation model, then it uses the newly created meshes to initialize a pose estimation and tracking model. Subsequently, the diffusion policy is conditioned on both the robot state and the continuously estimated object poses for the target objects. Crucially, because PRISM-DP uses a mesh generation model, it eliminates the need for manual mesh processing or creation, improving scalability and usability in open-set, real-world environments.
Through several experiments in both simulated and real-world settings, we demonstrate that PRISM-DP facilitates the learning of compact diffusion policies that achieve performance comparable to policies trained using ground-truth object pose information or manually generated meshes, while significantly surpassing image-based diffusion policies with similar parameter counts. To our knowledge, our work is the first to show that diffusion policies conditioned on pose-based observations can consistently outperform those conditioned on raw image observations in a real-world setting, suggesting exciting new directions for future research. Our implementation is publicly available, aiming to foster future research into scalable policy learning through open-set mesh generation and pose estimation.  [^1] https://github.com/Apollo-Lab-Yale/prism-dp
@@ 3 Technical Overview
In this section, we present a high-level overview of our approach, with further details on our specific implementation that instantiates this approach provided in the following section.
Our goal is to learn a compact and efficient receding horizon policy $\pi_{\theta}$ that, given a sequence of past observations $\mathbf{o}_{t-H_{o}:t}$, predicts a future action sequence $\mathbf{a}_{t:t+H_{p}}$, where $t$ denotes the current time step, $H_{o}$ is the observation horizon, and $H_{p}$ is the prediction horizon. Through this work, the $\pi_{\theta}$ policy is specifically learned via a diffusion policy strategy [1].
@@ 4.1 Prototype System
In this subsection, we provide implementation details for our prototype system that instantiates the approach outlined in §3.
For the $\mathrm{Segmentation}$ subroutine, the base model in our system is Segment Anything 2 (SAM2) [43]. Task-relevant objects are specified to SAM2 by clicking on the target objects within the sequence of RGB images on their first visible frames, $\mathbf{I}_{F_{i}}$. Such inputs are called “point prompts”. Our system also offers the option of segmenting using a YOLO11-Seg [44] model that is fine-tuned via SAM2 images. This option is particularly useful when segmentations must be performed across many similar images, as it reduces the need for tedious manual point prompts on all images. Instead, the user only needs to provide point prompts for a subset of images, after YOLO11-Seg can be fine-tuned to generalize segmentations over the specified target objects automatically.
@@ 4.2 Simulation Experiments
All models are implemented in PyTorch [48]. Training and evaluation of policies are performed on a workstation equipped with an AMD PRO 5975WX CPU, dual NVIDIA RTX 4090 GPUs, and 128GB of RAM. Each policy is trained for 200 epochs with batch size of 128, and evaluated over 250 rollouts.
@@
Table 1: Mean success rates (SR) and training time per epoch (TE) across five Robosuite tasks. TE is recorded in seconds. Evaluated methods include transformer-based (DP-T) and UNet-based (DP-U) diffusion policies trained with ground-truth object poses (GTP), raw RGB images (Img), and estimated 6D poses using ground-truth meshes (GTM). We also include larger image-based variants, DP-T-L Img and DP-U-L Img, that are 8 times and 11 times larger than their standard counterparts. Our proposed method with a transformer and a UNet is denoted as PRISM-DP-T and PRISM-DP-U.
@@ 4.3 Real-World Experiments
Experimental Settings.
To further validate the proposed approach, we conduct real-world experiments using a dual-arm robotic system (Fig. 6). The setup comprises two 7-DOF xArm7 manipulators, each mounted on a 1-DOF linear actuator. One arm is equipped with a RealSense D435i RGB–D camera for viewpoint control, while the other is fitted with a parallel gripper for manipulation. This configuration enables the policy to dynamically select viewpoints during task execution.
@@
Figure 6: Mean success rates (SR) and training time per epoch (TE) for all real-world tasks. The evaluated baselines and variants are the same as simulation experiments except for policies learned from ground-truth states since ground-truth states are no long accessible in real world.
@@ 5 Discussion
To our knowledge, this is the first work to demonstrate that diffusion policies conditioned on pose-based observations can consistently and significantly outperform those conditioned on raw image observations in real-world settings. While our current approach has several limitations that may hinder its immediate broad deployment (outlined below), this initial evidence alone underscores the potential of structured, pose-based observations as a promising direction for future research on diffusion policies or related imitation learning paradigms.
@@
Limitations. Our work has several limitations that suggest directions for future research. First, current mesh generation methods are either too computationally expensive or fail to produce meshes suitable for reliable pose estimation. The Meshy model [42] used in this work takes approximately one minute to generate a single mesh. Alternative methods we tested, such as [60, 38, 35], were unable to produce meshes of sufficient fidelity for downstream pose estimation. While generated meshes often achieve photorealistic appearance, their topology is far from ground-truth quality. For example, the generated can shown in Fig. 5 contains approximately $1.5\times 10^{5}$ triangular faces, whereas the corresponding ground-truth mesh has only 956. This excessive triangle count not only inflates memory usage but also increases rendering time during pose estimation and tracking, resulting in latency bottlenecks. Improving the speed, fidelity, and topological quality of mesh generation remains an important direction for future work. Promising approaches may include model distillation, mesh decimation, the use of geometric inductive biases, and incorporation of explicit topological constraints during mesh generation to produce compact yet accurate 3D models that better support efficient pose estimation.
@@
Second, our current pipeline and experiments use unit quaternions to represent the orientation of target objects. While quaternions provide a generic and widely applicable starting point, prior work suggests that alternative representations, such as 6D rotation representations [61], can offer improved learning stability and performance. Exploring more structured rotation parameterizations is an important avenue for future improvements.
Third, our current approach assumes that once a target object enters the camera’s field of view, it remains visible throughout the demonstration. This assumption was valid for the relatively uncluttered environments used in our experiments. However, in more realistic, cluttered settings where occlusions are common, this assumption would break down and impair tracking performance. Addressing this limitation will require future extensions that can robustly handle objects entering and exiting the field of view during task execution.
@@
Overall, while our approach demonstrates strong and promising initial performance, addressing these limitations will be crucial for scaling to more complex and dynamic real-world environments.

### v2

@@ 1 Introduction
Diffusion-based visuomotor policies generate robot motions by learning to denoise action-space trajectories conditioned on observations [1]. These observations are commonly sequences of raw RGB images, appealing due to their generality and ease of deployment, requiring just a camera. However, raw image sequences contain substantial amounts of irrelevant information, making it challenging for models to extract meaningful patterns without extensive computational resources, often necessitating hundreds of millions of parameters to fit demonstration data effectively.
@@
Recent work in learning-based 6D pose estimation and tracking, notably FoundationPose from Wen et al. [2], have introduced promising solutions for accurately estimating and tracking object poses from RGB-D video streams. In theory, these methods could enable compact pose-based diffusion policy training without external tracking setups. However, a significant limitation remains: FoundationPose and similar approaches depend on manually reconstructed or hand-crafted object meshes as input. Prior attempts to integrate FoundationPose into diffusion policy frameworks [3] have been constrained by the need for manual preprocessing of object meshes, hindering scalability in dynamic, open-set environments where object composition varies across tasks.
To address these challenges, we introduce Pose-tRacking via Image Segmentation and Mesh-generation Diffusion Policies (PRISM-DP), an approach that leverages segmentation, mesh generation, pose estimation, and pose tracking models to enable compact diffusion policy learning directly from the spatial poses of task-relevant objects. More specifically, PRISM-DP applies a segmentation model to isolate task-relevant objects in input images, it uses the resulting segmentation masks to automatically create 3D triangulated meshes for all target objects using a mesh generation model, then it uses the newly created meshes to initialize a pose estimation and tracking model. Subsequently, the diffusion policy is conditioned on both the robot state and the continuously estimated object poses for the target objects. Crucially, because PRISM-DP uses a mesh generation model, it eliminates the need for manual mesh processing or creation, improving scalability and usability in open-set, real-world environments.
Through several experiments in both simulated and real-world settings, we demonstrate that PRISM-DP facilitates the learning of compact diffusion policies that achieve performance comparable to policies trained using ground-truth object pose information or manually generated meshes, while significantly surpassing image-based diffusion policies with similar parameter counts. To our knowledge, our work is the first to show that diffusion policies conditioned on pose-based observations can consistently outperform those conditioned on raw image observations in a real-world setting, suggesting exciting new directions for future research. Our implementation is publicly available, aiming to foster future research into scalable policy learning through open-set mesh generation and pose estimation.  [^1] https://github.com/Apollo-Lab-Yale/prism
@@ 3 Technical Overview
Our goal is to learn a compact and efficient receding horizon policy $\pi_{\theta}$ that, given a sequence of past observations $\mathbf{o}_{t-H_{o}:t}$, predicts a future action sequence $\mathbf{a}_{t:t+H_{p}}$, where $t$ denotes the current time step, $H_{o}$ is the observation horizon, and $H_{p}$ is the prediction horizon.
@@ 4.1 Prototype System
In this subsection, we provide implementation details for our prototype system that instantiates the approach outlined in §3. For the $\mathrm{Segmentation}$ subroutine, the base model in our system is Segment Anything 2 (SAM2) [43]. Task-relevant objects are specified to SAM2 by clicking on the target objects within the sequence of RGB images on their first visible frames, $\mathbf{I}_{F_{i}}$. Such inputs are called “point prompts”.
@@
Importantly, our current system assumes a human-in-the-loop during task execution. Initially, the poses of all task-relevant objects are set to null values (all zeros). When a target object first appears in the field of view, the user pauses execution and provides point prompts to SAM2 through a GUI. FoundationPose then uses the resulting segmentation mask and corresponding mesh object (cached during training) to initialize the object’s pose via pose estimation. Execution then resumes, and FoundationPose continuously updates this object’s pose through tracking. This initialization process is performed separately for each target object.
@@ 4.2 Simulation Experiments
All models are implemented in PyTorch [47]. Training and evaluation of policies are performed on a workstation equipped with an AMD PRO 5975WX CPU, dual NVIDIA RTX 4090 GPUs, and 128GB of RAM. Each policy is trained for 200 epochs with batch size of 128, and evaluated over 250 rollouts to calculate the success rate by dividing the total number of evaluation rollouts with number of successful evaluation rollouts.
@@
Table 1: Success rates (SR) and training time per epoch (TE) across five Robosuite tasks. TE is recorded in seconds. Evaluated methods include Transformer-based (DP-T) and UNet-based (DP-U) diffusion policies trained with ground-truth object poses (GTP), raw RGB images (Img), and estimated 6D poses using ground-truth meshes (GTM). We also include larger image-based variants, DP-T-L Img and DP-U-L Img, that are 8 times and 11 times larger than their standard counterparts. Our proposed method with a Transformer and a UNet is denoted as PRISM-DP-T and PRISM-DP-U.
@@ 4.3 Real-World Experiments
Experimental Settings. To further validate the proposed approach, we conduct real-world experiments using a dual-arm robotic system (Fig. 6). The setup comprises two 7-DOF xArm7 manipulators, each mounted on a 1-DOF linear actuator. One arm is equipped with a RealSense D435i RGB–D camera for viewpoint control, while the other is fitted with a parallel gripper for manipulation. This configuration enables the policy to dynamically select viewpoints during task execution.
@@
Figure 6: Success rates (SR) and training time per epoch (TE) for all real-world tasks. The evaluated baselines and variants are the same as simulation experiments except for policies learned from ground-truth states since ground-truth states are no long accessible in real world.
@@ 5 Discussion
To our knowledge, this is the first work to demonstrate that diffusion policies conditioned on pose-based observations can consistently and significantly outperform those conditioned on raw image observations in real-world settings on suitable tasks. While our current approach has several limitations that may hinder its immediate broad deployment (outlined below), this initial evidence alone underscores the potential of structured, pose-based observations as a promising direction for future research on diffusion policies or related imitation learning paradigms.
@@
Limitations. Our work has several limitations that suggest directions for future research. First, current mesh generation methods are either too computationally expensive or fail to produce meshes suitable for reliable pose estimation. The Meshy model [42] used in this work, for instance, requires approximately one minute to generate a single mesh. Alternative methods we tested [59, 38, 35] were unable to produce meshes of sufficient fidelity for downstream pose estimation. Although generated meshes often achieve photorealistic appearance, their topology deviates substantially from ground-truth quality. For example, the generated can shown in Fig. 5 contains approximately $1.5\times 10^{5}$ triangular faces, whereas the corresponding ground-truth mesh has only 956. This excessive triangle count inflates memory usage and increases rendering time during pose estimation and tracking, resulting in latency bottlenecks. Improving the speed, fidelity, and topological quality of mesh generation remains an important direction for future work. Promising approaches may include model distillation, mesh decimation, the use of geometric inductive biases, and incorporation of explicit topological constraints to produce compact yet accurate 3D models that better support efficient pose estimation.
Our current system also assumes that once a target object enters the camera’s field of view, it remains visible throughout task execution. This assumption was not overly restricting in our relatively uncluttered experimental environments but would likely break down in more complex, dynamic settings where occlusions are common. Moreover, our current method for initializing object poses relies on a human-in-the-loop procedure: when a target object first appears, the user must manually pause execution and provide point prompts to SAM2 through a GUI to generate an initial segmentation mask for pose estimation. While effective, this procedure limits scalability to environments with many objects or frequent changes, and reducing this dependence on human input remains an important challenge.
Additionally, our approach is restricted to handling rigid-body motion. It cannot model nonrigid object deformations, such as folding a shirt or manipulating flexible materials–scenarios where, in theory, image-based observation may still succeed at recognizing necessary patterns.
@@
Another limitation stems from our use of unit quaternions to represent the orientation of target objects. While quaternions provide a broadly applicable starting point for reporting on initial results, prior work suggests that alternative rotation representations, such as continuous 6D parameterizations [60], can offer improved learning stability and performance. Exploring more structured orientation parameterizations is an important direction for future improvements.
@@
Overall, while our method demonstrates strong and promising initial performance, addressing these limitations will be critical for scaling to more complex, cluttered, and dynamic real-world environments.